OpenStudy (anonymous):

Find the linearization of f(x) = 6 cos x at x = −/4.

OpenStudy (anonymous):

"Find the linearization of..." just means find the equation of the tangent line at some particular x value. btw, at x = what???

OpenStudy (anonymous):

i said x= -1/4

OpenStudy (anonymous):

x = -1/4 ??? or is it \(\large x=-\frac{\pi}{4} \) ???

OpenStudy (anonymous):

do you have f'(x) = ???